OpenStudy (andriod09):

Betty and Alexander are standing in a big circle with 88 other people all facing the middle of the circle. Betty and Alexander are directly across from each other. The average width in the crowd is two feet. How far, approximitly are Betty and Alexander from each other?

OpenStudy (andriod09):

@hartnn

hartnn (hartnn):

The average width is 2 feet, so lets say all of them are approx 2 feet apart. Now , 88 people standing will have 88 gaps each of 2 feet. So, the circumference of circle will be 88*2 = 176 feet. ok with this ?

OpenStudy (andriod09):

there are 90 people in the circle though. 88+betty and alexander

OpenStudy (andriod09):

so it would be 180 feet.

hartnn (hartnn):

ohh!! sorry circumference = 90*2 = 180 feet yes. now can u find the diameter of circle ?

OpenStudy (andriod09):

dont remember. :[

hartnn (hartnn):

ok, Circumference = \(\pi\) *diameter. now whats the diameter ?

OpenStudy (andriod09):

90?

hartnn (hartnn):

since they both are directly across each other , the are on the end points of diameter so the distance between them is = diameter = 28.648 feet approx does this make sense ?

OpenStudy (andriod09):

not really. lol. i barly understand what 90/ \(\pi\) = 90/3.14 = 28.648 feet means, i cant see equations only drawing idk why though.

hartnn (hartnn):

\(\pi\) is an irrational number = 3.14 or 22/7 approx. and it is the ratio of circumference to diameter.

OpenStudy (andriod09):

ik what pi is. its all those slashes and parethesis and everything like that

hartnn (hartnn):

/ means division.

OpenStudy (andriod09):

\(\pi\)=i cannot see the equation so, just type it, dont use the equation button

OpenStudy (andriod09):

i can't see the \(\\) and what it means.

hartnn (hartnn):

ok, circumference = pi * diameter as circumference = 180 diameter = 180/ pi = 57.296 feet = distance between them

OpenStudy (andriod09):

wait, yo said the diameter is 28.648 feet approx

hartnn (hartnn):

that was typing error, when i saw all comments again , i found that circumference = 180 and not 90 so diameter = 57.296 feet.

OpenStudy (andriod09):

okay. thank you.

OpenStudy (andriod09):

i gave you a medal

hartnn (hartnn):

welcome :)